# Splat-Nav: Safe Real-Time Robot Navigation in Gaussian Splatting Maps

### v1

@@ III-E Supporting Hyperplanes
We need to also prove that the above constraint is a supporting hyperplane by showing that all feasible points when the constraint is equality is necessarily outside of the ellipsoid parametrized by $Q_{j}$. Recall that $x_{0}=\mu_{j}+\frac{\Delta_{j}}{k_{j}^{*}}$ is a point both on the surface of the ellipsoid and on the hyperplane. Therefore, all feasible points on the hyperplane can be expressed as $x=x_{0}+\delta$, where $\Delta_{j}^{T}Q_{j}\delta=0$. If the hyperplane supports the ellipsoid parametrized by $Q_{j}$, then necessarily $f(x)=(x-\mu_{j})^{T}Q_{j}(x-\mu_{j})\geq 1$. For all points on the hyperplane, the ellipsoid constraint evaluates to
$f_{j}(x_{0}+\delta)=\left(\frac{1+\epsilon}{k_{j}^{*}}\Delta_{j}+\delta\right)^{T}Q_{j}\left(\frac{1+\epsilon}{k_{j}^{*}}\Delta_{j}+\delta\right)\\ =\underbrace{\frac{(1+\epsilon)^{2}}{(k_{j}^{*})^{2}}\Delta_{j}^{T}Q_{j}\Delta_{j}}_{=(1+\epsilon)^{2}}+\underbrace{\delta^{T}Q_{j}\delta}_{\geq 0}+\underbrace{\frac{2(1+\epsilon)}{k_{j}^{*}}\Delta_{j}^{T}Q_{j}\delta}_{=0}>1.$
Therefore, the constraint in Proposition 2 is a supporting hyperplane. ∎
Figure 2: Our supporting hyperplane algorithm finds safe polytopes (green) in the presence of ellipsoidal obstacles (red) regardless of the safety of the query point. We also show that our algorithm is space-filling (and therefore non-conservative) by querying points on a grid (black dots) representing a robot with some radius. The resulting union of safe polytopes converges to the full extent of the collision-less space in the limit.
Corollary 4.
Based on Corollary 1, the polytopes will also be guaranteed safe regardless of the sampling scheme used to approximate $A_{j}$. Let $\hat{s}$ be the sampled $s$ that maximizes $K(s)$. We then use $\hat{s}$ to generate $\hat{Q}_{j}$ and $\hat{k}_{j}$. The point $\hat{x}_{0}$ calculated using the sampled values lies on the constraint hyperplane and is farther away from the ellipsoid than the true $x_{0}$ because the sampled $\hat{k}_{j}\leq k_{j}^{*}$.
This allows us to use the polytope generation with our parallelized sampling method in a safe way. We additionally note that the above constraints and polytopes apply to ellipsoidal robot bodies, as we never relied on the structure of $Q_{j}$ beyond it being symmetric and positive-definite. We believe that these findings related to ellipsoid-based environment representations may have an impact beyond the scope of safety, such as in graphics.
III-F Generating Paths
Our planning pipeline is amenable to sampling-based planners such as RRT and safe polytopic planners which generate safe corridors, described in the subsequent discussion.
III-F1 RRT
Splat-Nav can utilize an out-of-the-box RRT solver from OMPL [40], replacing the default collision checker with a custom checker that uses our formula from Corollary 3. This checker loads in the ellipsoids generated by GSplat, initializes a KD tree using the means, and then checks for collision for any sampled point within the RRT. We evaluate this method as a baseline compared to our proposed approach which plans through safe flight corridors.
III-F2 Safe Flight Corridors
Splat-Plan generates feasible motion plans using the safe flight corridor method from [27], leveraging Proposition 2 to generate safe polytopes. To do this, we create a 3D binary occupancy grid using only the means of the Gaussians. [^1] This may not result in a safe path, which we can handle but [27] cannot. We then use A* to generate a candidate path. Each point on the A* path almost always has sufficient free space in its neighborhood, and will not be stuck in corners since the path is fully connected from start to goal. However, there may be cases where the union of polytopes queried on the A* is disconnected, leading to an infeasible trajectory, which may happen when the A* path goes through walls of ellipsoids. To remedy this, we execute A* multiple times, running our collision test 3 for intersections and updating our occupancy grid accordingly. In this way, the final A* path is guaranteed to converge to a safe path. This fact, along with the fact that polytopes generated from safe points necessarily contain those points, guarantees the generation of a connected polytopic corridor. In Splat-Plan, we propose two approaches to planning:
Full Trajectory
In this approach, we first solve the planning problem for a full trajectory from the start to the goal. Using the iterated A* path, we generate a sequence of polytopes using Proposition 2. Subsequently, we find a minimal representation of the sequence of polytopes, since some constraints in the polytopes are redundant. Finally, we use the same method as Liu et al. [27] to find a smooth, dynamically feasible trajectory by solving a quadratic program. Since we are optimizing over the control points of a B-spline constrained to lie within the polytopes, we can guarantee that the paths are safe in the continuum, by the convex hull property of B-splines.
Local Trajectory
Although we find a minimal representation for the polytopes, these polytopes can still contain many facets, and there may be many polytopes in the corridor. We note that Liu et al. [27] sparsify their point cloud to reduce the number of constraints in their problem to achieve real-time performance. Instead, we take a model predictive control (MPC) perspective, planning a trajectory over a receding time horizon and replanning before the robot reaches the end of this partial trajectory. Using the same set of polytopes (which are quick to compute) as the Full Trajectory, our MPC problem uses the current safe polytope and plans a path to the A* waypoint in the next polytope. This greatly reduces the number of constraints in the trajectory optimization problem, allowing us to achieve real-time performance. Because this one-step look-ahead method still uses the same set of polytopes as the Full Trajectory, the solved trajectories will still be guaranteed safe, at the expense of sub-optimality.
Remark 1.
An arbitrary, continuous sequence of points (e.g., a straight line from start to goal) may not generate a connected sequence of polytopes. This may occur when a particular point offers no safe polytope (i.e., a point is immediately surrounded on all sides by ellipsoids) or the sequence of points biases the polytopes into a corner so that polytopes exist on both sides of the corner, but not in the corner itself. While there may exist ways to carefully account for these edge-cases when building a trajectory, we leave this to future work.
IV Pose Estimation
We present our pose estimation module, Splat-Loc, for localizing a robot in a Gaussian Splatting representation of its environment. We assume that the robot has an onboard RGB-D camera with known camera intrinsics. We note that prior knowledge of an estimate of the robot’s pose (such as its neighborhood) may not be available in many practical settings. As such, we design a pose estimation module that can perform global localization (i.e., pose initialization) and can utilizes prior knowledge of the pose when available to speed up operation (i.e., a recursive estimation algorithm). We assume that the robot has an onboard camera with known camera intrinsics, capturing images of its environment, which we utilize in estimating the robot’s pose.
IV-A Global Initialization
When a good initial guess of the robot’s pose is unavailable, we execute a global pose estimation procedure. We first use the RGB-D image and the known camera intrinsics to generate the point cloud (in the camera frame). In addition, we generate a point cloud of the scene from the GSplat using the ellipsoid means $\{\mu_{j}\}_{j=1}^{N}$. This then becomes a point-cloud registration task, which we can leverage existing tool to solve. The point-cloud registration problem is given by:
$T^{*}=\operatorname*{arg\,min}_{T\in\mathrm{SE}(3)}\sum_{(p,q)\in\mathcal{C}}\left\|p-Tq\right\|_{2}^{2},$ (3)
where the transformation matrix $T\in\mathrm{SE}(3)$} comprises of a translation component $t\in\mathbb{R}^{3}$} and a rotation matrix $R\in\mathrm{SO}(3)$} and C{C} denotes the set of correspondences, associating the points between the two point clouds (where $p$ is in the map cloud and $q$ is in the sensor cloud). Given a known set of correspondences, we can compute the optimal solution of (3) using Umeyama’s method [47].
In practice, we do not have prior knowledge of the set of correspondences $\mathcal{C}$ between the two point clouds. To address this challenge, we apply standard techniques in feature-based global point-cloud registration. We begin by computing $33$-dimensional Fast Point Feature Histograms (FPFH) descriptors [38] for each point in the point-cloud, encoding the local geometric properties of each point. Prior work has shown that visual attributes can play an important role in improving the convergence speed of point-cloud registration algorithms [34], something that FPFH does not do. To solve this, we augment the FPFH feature descriptor of a given point with its RGB color. We then identify putative sets of correspondences using a nearest-neighbor query based on the augmented FPFH descriptors then utilize Random Sample Consensus (RANSAC) to iteratively remove identify outliers to get the final $\mathcal{C}$. The RANSAC convergence criterion is based on the distance between the aligned point clouds and the length of a pair of edges defined by the set of correspondences.
Remark 2.
We assume the ellipsoids in $\mathcal{G}$ are small enough (in a relative sense) such that the collection of Gaussians can be approximated by a point cloud. This assumption is automatically valid when camera viewpoints are far enough away from the scene that the point cloud projected onto the image plane is not sparse. Even if this is not the case, there are two ways in which sparsity can come into play: Gaussian-to-point sparsity and representation sparsity. If there exists large Gaussians in the scene or field of view, these can be subdivided into smaller children Gaussians without any degradation of quality to the representation. This is identical to sampling large Gaussians to add to the point cloud instead of simply taking the center. If there are no Gaussians in a particular area, trying to draw any results like a pose estimate is ill-posed because the representation lacks any information there.
IV-B Recursive Pose Estimation
The recursive pose estimators initializes itself from either a known initial pose or the the global estimate. It then allows the robot to efficiently track its pose as it moves. Our recursive pose estimator requires only an RGB image (no depth).
We first render a set of 2D images of the GSplat map within a local neighborhood of the previously known pose. We then identify keypoints and the associated descriptors for points in these map images as well as the robot’s RGB image using the encoder-decoder-based method SuperPoint [6]. Next, we match features between both images using the transformer-based method LightGlue [26]. [^2] Other methods for feature detection and matching can be utilized, such as Scale-Invariant Feature Transform (SIFT) [5] and Oriented FAST and Rotated BRIEF (ORB) [37]. We again utilize RANSAC to remove outliers from the set of matched features to yield the final correspondence set $\mathcal{C}$}. Finally, we take the map points from $\mathcal{C}$ and project them into the 3D scene using the perspective projection model, camera intrinsics, and the rendered depth from GSplat.
$T^{*}=\operatorname*{arg\,min}_{T\in\mathrm{SE}(3)}\sum_{(p,w)\in\mathcal{C}}d(\Pi(T,p),w),$ (4)
where $T\in\mathrm{SE}(3)$} denotes the camera extrinsic parameters, $\mathcal{C}$ denotes the set of correspondences between points in the environment $p$ and points in the observed RGB image $w$, $\Pi:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$} represents the perspective projection model of point $p$ using camera extrinsics $T$, and $d:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}$} denotes the reprojection error. One common choice for $d$ is the $\ell_{2}$-squared-norm, which we also utilize in this work. This yields a non-linear least-squares optimization problem that can be solved using variants of the Levenberg-Marquardt and Gauss-Newton optimization algorithms. We note that various methods have been developed to solve the PnnP problem such as the EPnnP [25], DLS [14], and SQPnnP [44] methods.
Remark 3.
In many robotics problems, an estimate of the covariance associated with the estimated pose of the robot is important for downstream decision-making and motion planning. In this work, we compute the Hessian of (4) at $T^{*}$ to estimate the covariance associated with the estimated pose of the robot.
Remark 4.
Our method can be directly extended to the batch-estimation case where we solve a bundle adjustment problem over multiple time steps.
V Results
We examine the performance of each component of our navigation and pose estimation pipeline and perform ablative studies, comparing our algorithms to existing methods. We evaluate our proposed pose estimator in one synthetic scene Stonehenge and two real scenes (trained from real RGB images) Statues and Flightroom, shown in Fig. 3.
V-A Safe Planning
We evaluate Splat-Plan, our polytopic safe corridor method, on 100 start and goal locations distributed across a circle in the three environments. We also compare it against baselines and prior work. Namely, we benchmark against: (1) an A* planner using the GSplat means as an occupancy grid, (2) an RRT* planner using the collision test 3, and (3) a gradient-based planner built for neural-network NeRFs, Nerf-Nav [1].
We also present several ablations and extensions of Splat-Plan. First, we investigate the properties of a more probabilistically grounded perspective on the ellipsoidal GSplat representation. Specifically, we compare using just the ellipsoid represented by the GSplat covariances (Splat-Plan Basic) to scaling the ellipsoids by a factor $\beta_{j}\alpha_{j}$, where $\beta_{j}$ is related to the confidence interval of the Gaussian and $\alpha_{j}$ is the opacity of the Gaussian. This scaling therefore represents the joint probability of an object existing in an ellipsoidal region of space and that of the object being solid. For these experiments, we take $\beta_{j}=2$.
We also benchmark our claims between the Full Trajectory and Local Trajectory methods (Iterative A* following), an ablation of Full Trajectory that does not iterate on A* (Non-iterative A*), and a Local Trajectory method that naively uses a straight line path to the goal to create polytopes instead of A*. Then present hardware experiments onboard a drone in the Flightroom environment. For all tests, we parametrize the robot body as a sphere with radius $r=0.03$, except for the hardware tests, where $r=0.28$ m.
Visually, the paths generated by Splat-Plan are smooth, safe, and non-conservative (Fig. 3). This fact is validated in Fig. 4, where Splat-Plan’s trajectories (probabilistic and basic variants) are safe (distances greater than 0) with respect to the ground-truth mesh in a vast majority of trials. Moreover, these trajectories are non-conservative due to small path lengths and variances. We also see that the probabilistic variant is slightly safer and has marginally smaller path lengths. Using the GSplat means and a gradient-based planner has no performance guarantees and therefore suffer in safety. The RRT* also suffers due to tradeoff between resolution of the collision checking and the lack of scalability in sampling random points.
@@ V-A Safe Planning
Although the Full Trajectory method presented in Fig. 4 is generalizable, we wanted to test the limits of the safe corridor generation by benchmarking the Local Trajectory method and removing components of the Full Trajectory. We see that all methods are safe (Fig. 5) due to the robust guarantees of Proposition 2. The Local Trajectory method (Iterative A* Following) always succeeds in finding feasible trajectories to the goal, validating our performance claim. The sub-optimality of the path lengths are also apparent. Without relying on the iterative A* procedure (Non-Iterative A*), the feasibility and connectivity of trajectories falls to 55% due to the “corner” phenomenon (Remark 1). Interestingly, removing A* entirely and sampling corridor query points on the line between the current and goal location (Line Following) is surprisingly performant, able to navigate to the goal 58% of the time. Finally, we executed Splat-Plan (probabilistic variant) on the real Flightroom environment onboard a drone (r = $0.28$ cm). Visually and qualitatively, the executed trajectory is safe, which we demonstrate in Figure 6.
Figure 4: A comparison of trajectories generated by our method and 3 other baselines (A*, RRT*, and NeRF-Nav [1]) for the same 100 start and end locations for each scene (Stonehenge, Statues, and Flightroom). We evaluate the safety and conservativeness of trajectories based on distance to the scene mesh and the length of the path, respectively. Each dot represents a particular trajectory’s minimum distance to the mesh and its path length, and horizontal markings indicate the max, mean, and min over all trajectories. By using a probabilistic weighting on the scaling of the ellipsoids in the representation, our proposed method is able to be safe (min. distance greater than 0) and is not conservative (small path length and variance).
@@
Figure 5: Extensions and ablations of Splat-Plan evaluated on the same configurations as Figure 4 on Stonehenge. We compare three methods: iterative A* following, non-iterative A* following, and line following. All the methods are still safe (distances greater than 0) because they use polytopic corridors. Iterative A* has a 100% success rate by nature of the interaction between the safe polytopes and a guaranteed safe A* path. Non-iterative A* could not find a solution 45% of the time due to the inability to find a connected union of polytopes generated by the A*. Finally, line following produced infeasible programs 18% of the time, and timed out getting to the goal (i.e. ”stuck”) 24% of the time.
Figure 6: (top) Deployment of our safe planner, Splat-Plan, on a drone. (bottom-left) Visualization of the executed trajectory (green) and the commanded trajectory (black outline). A straight line path from start to goal is not safe. (bottom-right) Distance of the drone to the closest obstacle, taking into account the extent of the drone. Note that the executed trajectory lies above the red line and is therefore safe.
V-B Pose Estimation
We compare our proposed pose estimator to other pose estimation methods, including Point-to-plane Iterative Closest Point (ICP) [4] and Colored-ICP [34]. In addition, we examine two variants of our pose estimator: Splat-Loc-GLue, which utilizes LightGlue for feature matching; and Splat-Loc-SIFT, which utilizes SIFT for feature matching. In each scene, we run 10 trials (of $100$ frames each) of each pose estimation algorithm. We evaluate the rotation and translation errors with respect to the ground-truth pose as well as the computation time per frame and the success rate (a success means it generated a solution, regardless of quality).
The performance of pose estimation algorithms often depends on the error associated with initial estimate of the pose. As such, we seek to examine these algorithms across settings with different errors in the initial estimate of the pose. In this study, we do not utilize the initialization procedure described in Section IV, rather we assume an initial estimate of the pose is available. We generate the initial estimate by taking the ground truth pose then applying a rotation $\delta_{R}$ about a random axis and the translation $\delta_{t}$ in a random direction.
V-B1 Pose Initialization of Splat-Loc
We evaluate the global point-cloud alignment initialization procedure as well as the performance of Splat-Loc-Glue in the Statues scene (a real-world scene) and Stonehenge (a synthetic scene), with the results in Table I. We see that the success rate is 80% in both scenes and it runs at approximately 20-30 Hz. When it does succeed, the pose errors are on the order of 0.3 deg and 1 cm. Although these estimates are relatively good, we note that these are less accurate than the recursive pose estimates discussed later.
TABLE I: Performance of the pose initialization module in Splat-Loc
Scene Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$)
Statues $3.683\mathrm{e}^{-1}\pm 7.207\mathrm{e}^{-1}$ $1.332\mathrm{e}^{-2}\pm 2.329\mathrm{e}^{-2}$ $3.387\mathrm{e}^{-2}\pm 5.696\mathrm{e}^{-4}$ $80$
Stonehenge $2.138\mathrm{e}^{-1}\pm 2.068\mathrm{e}^{-1}$ $3.026\mathrm{e}^{-3}\pm 2.274\mathrm{e}^{-3}$ $4.572\mathrm{e}^{-2}\pm 1.697\mathrm{e}^{-3}$ $80$
V-B2 Recursive Pose Estimation
Table II shows our performance metrics in the Statues scene with with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$. First, we see that only the Colored-ICP algorithm fails to achieve a $100\%$ success rate. Second, our algorithms Splat-Loc-Glue and Splat-Loc-SIFT yield higher-accuracy and more consistent solutions, with mean rotation and translation errors less than 0.1 deg and 7 mm, respectively, and lower standard deviations. In these tests, Splat-Loc-Glue performs the best on each metric with a computation time of about 35 ms. Figure 7 shows a run of the algorithms where ICP and Colored-ICP achieve relatively lower errors.
TABLE II: Comparison of pose estimation algorithms in the Statues scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$) ICP [4] $8.416\mathrm{e}^{1}\pm 5.578\mathrm{e}^{1}$ $3.829\mathrm{e}^{0}\pm 8.507\mathrm{e}^{0}$ $8.595\mathrm{e}^{-2}\pm 4.398\mathrm{e}^{-2}$ $100$ Colored-ICP [34] $3.020\mathrm{e}^{1}\pm 4.573\mathrm{e}^{1}$ $3.455\mathrm{e}^{-1}\pm 5.102\mathrm{e}^{-1}$ $6.270\mathrm{e}^{-2}\pm 2.716\mathrm{e}^{-2}$ $80$ Splat-Loc-SIFT (ours) $8.314\mathrm{e}^{-2}\pm 9.367\mathrm{e}^{-3}$ $5.720\mathrm{e}^{-3}\pm 7.619\mathrm{e}^{-4}$ $6.557\mathrm{e}^{-2}\pm 7.158\mathrm{e}^{-4}$ $100$ Splat-Loc-Glue (ours) $\bm{5.726\mathrm{e}^{-2}\pm 4.815\mathrm{e}^{-3}}$ $\bm{3.129\mathrm{e}^{-3}\pm 3.232\mathrm{e}^{-4}}$ $\bm{3.412\mathrm{e}^{-2}\pm 1.992\mathrm{e}^{-3}}$ $100$
Figure 7: Statues Scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$—Pose error and computation time per frame achieved by each pose estimation algorithm in a specific instance of the problem where ICP and Colored-ICP perform well.
Next, Table III shows results for the Statues scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$. Note that the success rates of Colored-ICP and Splat-Loc-SIFT decreases sharply, suggesting that these methods require a good initial estimate of the pose of the robot. However, when they do succeed, these algorithms achieve low rotation and translation errors, with Splat-Loc-SIFT achieving a much-higher accuracy compared to Colored-ICP. In contrast, the Splat-Loc-Glue algorithm achieves a perfect success rate, along with the lowest rotation and translation errors and computation time. The results suggest the greater robustness of Splat-Loc-Glue compared to Splat-Loc-SIFT. Figure 8 shows the performance of each algorithm in an instance of the problem where all methods are successful.
TABLE III: Comparison of pose estimation algorithms in the Statues scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$) ICP [4] $9.897\mathrm{e}^{1}\pm 3.051\mathrm{e}^{1}$ $1.401\mathrm{e}^{0}\pm 2.297\mathrm{e}^{-1}$ $4.354\mathrm{e}^{-2}\pm 5.071\mathrm{e}^{-2}$ $100$ Colored-ICP [34] $2.419\mathrm{e}^{0}\pm 1.609\mathrm{e}^{0}$ $2.183\mathrm{e}^{-1}\pm 2.016\mathrm{e}^{-1}$ $7.029\mathrm{e}^{-2}\pm 2.435\mathrm{e}^{-2}$ 2020 Splat-Loc-SIFT (ours) $7.844\mathrm{e}^{-2}\pm 5.953\mathrm{e}^{-3}$ $6.275\mathrm{e}^{-3}\pm 8.400\mathrm{e}^{-4}$ $7.312\mathrm{e}^{-2}\pm 1.498\mathrm{e}^{-3}$ $30$ Splat-Loc-Glue (ours) $\bm{6.139\mathrm{e}^{-2}\pm 5.994\mathrm{e}^{-3}}$ $\bm{3.671\mathrm{e}^{-3}\pm 6.535\mathrm{e}^{-4}}$ $\bm{4.278\mathrm{e}^{-2}\pm 1.498\mathrm{e}^{-3}}$ $100$
Figure 8: Statues Scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$—Pose error and computation time per frame achieved by each pose estimation algorithm in an instance where Colored-ICP and Splat-Loc-SIFT are successful.
We present additional results on the Flightroom scene in the Appendix.
VI Conclusion
We introduce an efficient navigation pipeline termed Splat-Nav for robots operating in Gaussian Splatting environments. Splat-Nav consists of a guaranteed-safe planning module Splat-Plan, which allows for real-time planning (5 Hz) by leveraging the ellipsoidal representation inherent in GSplats for efficient collision-checking and safe corridor generation, facilitating real-time online replanning. Moreover, our proposed pose estimation module Splat-Loc computes high-accuracy pose estimates faster (20 Hz) and more reliably compared to existing pose estimation algorithms for radiance fields, such as NeRFs. We present extensive simulation results, highlighting the effectiveness of Splat-Nav. In future work, we seek to more deeply explore utilization of Splat-Nav across a diversity of hardware platforms and real-life environments.
Appendix
A Planning Computation Times
In addition to outperforming other existing methods in safety and path length of computed paths (Fig. 4,5), our Splat-Plan method achieves faster computation times compared to the point-based motion planners RRT* and NeRF-Nav; generates smooth trajectories unlike the RRT*, NeRF-Nav, and A*; and detects infeasible planning problems. We remind the reader of the following baselines: RRT*, A*, and polytopic A*. RRT* (implemented in OMPL) replaces the default collision checker with our ellipsoid-to-ellipsoid collision test in Proposition 1. The A* methods use the means of the Gaussians in GSplat to create a binary occupancy grid. A* simply plans a shortest path through the binary grid. Polytopic A* goes one step further, constructing safe polytopes through Proposition 2 on the A* waypoints, from which it computes a B-spline trajectory passing through all the polytopes. Splat-Plan is a further refinement of the Polytopic A* algorithm. Instead of a single instance of A*, we iterate between querying A* and updating the binary occupancy grid at the A* waypoints that are unsafe via the collision test in Proposition 1. Typically, the Splat-Plan method requires only a handful of A* iterations (<5{<5}), and the generation of polytopes for approximately $100$ A* waypoints runs at approximately $100$ Hz.
Splat-Plan is much faster than RRT* and NeRF-Nav [1] due to our method’s inherent sampling efficiency, as noted in Table IV. The faster computation time in Splat-Plan results primarily from the fast safe polytope generation procedure. Because both of these baselines also test/measure safety point-wise, they have finite resolution. Therefore, these baselines may not dynamically feasible for some robotic tasks, and may be unsafe in certain regions of the computed trajectory, if the resolution is insufficient.

### v2

@@ Appendix E Proof of Proposition 2
We need to also prove that the above constraint is a supporting hyperplane by showing that all feasible points when the constraint is equality is necessarily outside of the ellipsoid parametrized by $Q_{j}$. Recall that $x_{0}=\mu_{j}+\frac{\Delta_{j}}{k_{j}}$ is a point both on the surface of the ellipsoid and on the hyperplane. Therefore, all feasible points on the hyperplane can be expressed as $x=x_{0}+\delta$, where $\Delta_{j}^{T}Q_{j}\delta=0$. If the hyperplane supports the ellipsoid parametrized by $Q_{j}$, then necessarily $f(x)=(x-\mu_{j})^{T}Q_{j}(x-\mu_{j})\geq 1$. For all points on the hyperplane, the ellipsoid constraint evaluates to
$f_{j}(x_{0}+\delta)=\left(\frac{1+\epsilon}{k_{j}}\Delta_{j}+\delta\right)^{T}Q_{j}\left(\frac{1+\epsilon}{k_{j}}\Delta_{j}+\delta\right)\\ =\underbrace{\frac{(1+\epsilon)^{2}}{k_{j}^{2}}\Delta_{j}^{T}Q_{j}\Delta_{j}}_{=(1+\epsilon)^{2}}+\underbrace{\delta^{T}Q_{j}\delta}_{\geq 0}+\underbrace{\frac{2(1+\epsilon)}{k_{j}}\Delta_{j}^{T}Q_{j}\delta}_{=0}>1.$
Hence, the constraint in Proposition 2 is a supporting hyperplane.
Appendix F Extended Planning Results
We remind the reader of the following baselines: RRT*, A*, and polytopic A*. In RRT* (implemented in OMPL), we replace the default collision checker with our ellipsoid-to-ellipsoid collision test from Corollary 1. The A* methods use the means of the Gaussians in GSplat to create a binary occupancy grid. A* simply plans a shortest path through the binary grid. Polytopic A* goes one step further, constructing safe polytopes through Proposition 2 on the A* waypoints, from which it computes a B-spline trajectory passing through all the polytopes. Splat-Plan is a further refinement of the Polytopic A* algorithm. Instead of a single instance of A*, we iterate between querying A* and updating the binary occupancy grid at the A* waypoints that are unsafe via the collision test in Proposition 1. Typically, the Splat-Plan method requires only a handful of A* iterations (<5{<5}), and the generation of polytopes for about $100$ A* waypoints runs at about 100 Hz.
A Ablations
@@
We test the limits of the safe corridor generation by benchmarking the Splat-Plan MPC method and the Splat-Plan Global method. We see that all methods are safe (Fig. 13) due to the robust guarantees of Proposition 2. The Splat-Plan MPC (Iterative A* Following) always succeeds in finding feasible trajectories to the goal, validating our performance claim. The sub-optimality of the path lengths are also apparent. Without relying on the iterative A* procedure (Non-Iterative A*), the feasibility and connectivity of trajectories falls to 55% due to the “corner” phenomenon (Remark 1). Interestingly, removing A* entirely and sampling corridor query points on the line between the current and goal location (Line Following) is surprisingly performant, able to navigate to the goal 58% of the time.
@@
Figure 13: Extensions and ablations of Splat-Plan evaluated on the same configurations in Stonehenge. We compare three methods: iterative A* following, non-iterative A* following, and line following. All the methods are still safe (distances greater than 0) because they use polytopic corridors. Iterative A* has a 100% success rate by nature of the interaction between the safe polytopes and a guaranteed safe A* path. Non-iterative A* could not find a solution 45% of the time due to the inability to find a connected union of polytopes generated by the A*. Finally, line following produced infeasible programs 18% of the time, and timed out getting to the goal (i.e. ”stuck”) 24% of the time.
B Splat-Plan Times
In addition to outperforming other existing methods in safety and path length of computed paths (Fig. 13), our Splat-Plan method achieves faster computation times compared to the point-based motion planners RRT* and NeRF-Nav; generates smooth trajectories unlike the RRT*, NeRF-Nav, and A*; and detects infeasible planning problems.
TABLE III: Comparison of Full Trajectory Computation Times
[…]
Since Splat-Plan requires the use of A* to initialize a path to generate the safety polytope corridors, it will necessarily be slower than these A* baselines. As Table III shows, Splat-Plan is much faster than RRT* and NeRF-Nav [10] due to our method’s inherent sampling efficiency. The faster computation time in Splat-Plan results primarily from the fast safe polytope generation procedure. The dominating cost in Splat-Plan Global is the calculation of the Bézier spline due to the number of polytopes in the constraints, with many facets per polytope. We solved the QP using CVXPY with the Clarabel solver. We note that the computation time associated with the Bézier spline computation procedure can be reduced using embedded solvers or faster programming languages.
C Discussion of Safety